Two opposite point charges of magnitude 2.100e-6 C are located near the origin. The positive charge is at (1.50, 1.50) mm and the negative charge is at (−1.50, −1.50) mm. What is the magnitude of the dipole moment vector for this pair of charges?

Thanks!
Answer #1

dipole moment vector is given by

p = q*d

q = 2.1*10^-6 C

d = sqrt(3^2 + 3^2)

d = 4.24 mm

p = 2.1*10^-6*4.24*10^-3

p = 8.9*10^-9 C-m